# David Marín111david.marin”AT”uab.cat

### v2

@@ Abstract
Following [6, 7] we define a fake saddle as a singular point having exactly two separatrices and two hyperbolic sectors, see Figure 1. Another way to think about a fake saddle is that the corresponding vector field can be put in a form that is like a degenerate flow box, i.e. near the singularity the phase portrait consists of parallel fibers, all but one of which have no singular points, and the singular fiber (the union of the singular point with its two separatrices) has a semi-stable equilibrium point, see [2]. The local phase portrait of a singularity with non-zero linear part (for the nilpotent case see for instance [1, Theorem 3.5]) is well known and prevents the singular point from being a fake saddle. The first non-trivial (and generic) case of false saddle occurs when the second-order jet is non-zero. Our goal is to characterize such generic fake saddles and also to study their transition maps between two transverse sections to the singular fiber in order to know its attractive or repulsive character. The techniques we use allow to give the first term of an asymptotic expansion which is uniform with respect to any parameters of the fake saddle.
@@
There exist nilpotent singular points that are not fake saddles for which $y=0$ is a separatrix having a transition map defined only on one side $\pm y>0$. For instance, the vector field
$(\lambda y+x^{2}+y^{2})\partial_{x}-\frac{xy}{\lambda}\partial_{y},\quad\lambda>0,$
with Darboux first integral $|y|^{2\lambda}\left(x^{2}+\frac{\lambda}{\lambda+1}y^{2}+\frac{2\lambda^{2}}{1+2\lambda}y\right)$, has a well-defined transition map along $y=0$ only for $y>0$. The singularity at the origin has four characteristic orbits (or separatrices) $\{\pm x>0,y=0\}$, $\{\pm x>0,x^{2}+\frac{\lambda}{\lambda+1}y^{2}+\frac{2\lambda^{2}}{1+2\lambda}y=0\}$, separating four hyperbolic sectors. $\square$
This work is mainly motivated by the paper [2] which provides a normal form for generic fake saddles that we recall in the sequel. First we can locate the fake saddle at the origin $(0,0)$. Secondly, if the vector field is at least of class $\mathscr{C}^{3}$ there exists a local diffeomorphism that rectifies the germ of singular fiber, so that it is contained in the line $y=0$ in a certain coordinate system $(x,y)$, see [2, Lemma 2.1]. Thanks to the hypothesis that the second-order jet of the fake saddle is non-zero, after a suitable rescaling of the coordinates we can assume that the fake saddle is defined by a system of differential equations
@@
where $a,b,c\in\mathbb{R}$. $f_{1}(0,0)=f_{2}(0,0)=1$. This normal form motivates us to consider a smooth family of planar vector fields $\{X_{\mu}\}_{\mu\in\Omega\subset\mathbb{R}^{n}}$ having the form
@@
where $f_{1}(x,y;\mu)$, $f_{2}(x,y;\mu)$, $g_{1}(x,y;\mu)$ and $g_{2}(y;\mu)$ are $\mathscr{C}^{K}$-functions fulfilling $f_{1}(0,0;\mu)=f_{2}(0,0;\mu)=1$. We define the following invariants
$a(\mu),\quad b(\mu)=g_{2}(0;\mu),\quad c(\mu)=g_{1}(0,0;\mu)$ (2)
of the family (1) and consider the associated value $d(\mu):=4(1-c(\mu))-(a(\mu)-b(\mu))^{2}$ which will play a key role in the sequel.
@@
Theorem 2.
If the invariants $(a,b,c)$ given in (2) of a vector field $X$ in the family (1) do not belong to $\{d=0\}\cap\{a^{2}-b^{2}=4\}$ then the origin is a fake saddle if and only if, either $d>0$, or $c=1$ and $a=b$. In both situations after blowing up the origin we have a single singular point on the exceptional divisor, which is a hyperbolic saddle of hyperbolicity ratio 1c1-c in the first one and a semi-hyperbolic saddle in the second one.
Proof.
We have that
@@
with $Q(0,v)=-v^{2}+(b-a)v+c-1$. The point $(u,v)=(0,0)$ is always a singular point of $Y$. If $d<0$ then $Y$ has two other singular points on the exceptional divisor $u=0$ with at least one non-zero eigenvalue. If $(a,b,c)\in\{d=0\}\setminus(\{c=1,a=b\}\cup\{a^{2}-b^{2}=4\})$ then $Y$ has another double singular point at (u,v)=(0,(ba)/2)(u,v)=(0,(b-a)/2) with a non-zero eigenvalue. In these situations the transition map $\Pi_{\alpha}^{\omega}:\{x=\alpha\}\to\{x=\omega\}$ is not well defined so that the origin is not a fake saddle of $X$. The remaining assertions are easy to check. ∎
The hypothesis $(a,b,c)\notin\{d=0\}\cap\{a^{2}-b^{2}=4\}$ in Theorem 2 (which is always verified when $a=0$) can not be removed as the following example shows.
For each integer $n\geq 3$ we consider the vector field
$X_{n}=(x+y)^{2}\partial_{x}+y^{n}\partial_{y}$
On the other hand, the blowup of the vector field $X_{4}$
Moreover, since $a=b=0$, according to Theorem 4 below, the derivative at $v=0$ of the transition map $\Pi_{\alpha}^{\omega}(v)$ of $Y_{1}$ is
Let $\Omega$ be an open subset of $\mathbb{R}^{n}$ and let $W$ any subset of $\Omega$. Recall (cf. [4, Definition 1.2]) that a $\mathscr{C}^{K}$-function $f(s;\mu)$ defined in the intersection of $(0,+\infty)\times\Omega$ with an open neighborhood of $\{0\}\times\Omega\subset\mathbb{R}^{n+1}$ belongs to the flat class $\mathcal{F}_{L}^{K}(W)$ if for each $\mu_{0}\in W$ and $\nu=(\nu_{0},\ldots,\nu_{n})\in\mathbb{N}^{n+1}$ with $|\nu|=\nu_{0}+\cdots+\nu_{n}\leq K$ there exists $\varepsilon>0$ such that $|\partial_{s}^{\nu_{0}}\partial_{\mu_{1}}^{\nu_{1}}\cdots\partial_{\mu_{n}}^{\nu_{n}}f(s;\mu)|\leq Cs^{L-\nu_{0}}$ for every $s\in(0,\varepsilon)$ and $\mu\in\mathbb{B}_{\varepsilon}(\mu_{0})\cap\Omega$.
We consider the transverse sections $\Sigma_{\star}=\{x=\star\}$ to $y=0$, for $\star\in\{\alpha,\omega\}$, where $\alpha<0<\omega$. We assume that $f_{1}(x,0)>0$ for all $x\in[\alpha,\omega]$ and we consider the transition map $\Pi_{\alpha}^{\omega}:\Sigma_{\alpha}\to\Sigma_{\omega}$. Our second result extends [2, Theorem B], which only treats the case $a=b=0$ when the transition map is restricted to $y>0$.
Theorem 4.
Assume that $\mu_{0}\in\Omega$ and the invariants (2) of the vector field $X_{\mu_{0}}$ in (1) belong to $\{d=4(1-c)-(a-b)^{2}>0\}$. If $\mu\approx\mu_{0}$ then the origin is a fake saddle of $X_{\mu}$ and its transition map $\Pi_{\alpha}^{\omega}:\Sigma_{\alpha}\to\Sigma_{\omega}$ satisfies $\Pi_{\alpha}^{\omega}(y;\mu)=e^{\gamma_{\pm}(\mu)}y+\mathcal{F}_{1+\epsilon}^{K}(\{d>0\})$ on $\pm y\geq 0$, where $\epsilon>0$ and
Clearly the sign of $\gamma_{\pm}(\mu)$ determines the stability of the singular fiber $y=0$ on the side $\pm y>0$. In the case $a=b=0$ the two values $\gamma_{\pm}(\mu)$ coincide with the principal value of the integral. Otherwise the stability of the two sides $\pm y>0$ can be different, see Example Fake saddles and their transition maps 00footnotetext: 2020 AMS Subject Classification: 34D10, 34D20; Secondary: 34C05. 00footnotetext: Key words and phrases: singularities, Poincaré and Dulac transition maps, asymptotic expansion, uniform flatness, stability 00footnotetext: This work has been partially funded by the Ministry of Science, Innovation and Universities of Spain through the grant PID2021-125625NB-I00 and by the Agency for Management of University and Research Grants of Catalonia through the grant 2021SGR01015. . $\square$
Using that $f_{1}(0,0)=1$ and $g_{1}(0,0)=c$ we can compute the previous principal value as
$\mathrm{PV}\int_{\alpha}^{\omega}\frac{g_{1}(x,0)}{xf_{1}(x,0)}dx:=\lim\limits_{\varepsilon\to 0^{+}}\int_{(\alpha,\omega)\setminus(-\varepsilon,\varepsilon)}\frac{g_{1}(x,0)}{xf_{1}(x,0)}dx=c\log\left|\frac{\omega}{\alpha}\right|+\int_{\alpha}^{\omega}\left(\frac{g_{1}(x,0)}{f_{1}(x,0)}-c\right)\frac{dx}{x}.$
$\square$
Proof.
@@
Notice that $\lambda:=\lambda_{+}=1-c>0$ and and $\lambda_{-}=\frac{1}{\lambda_{+}}$. We have that
@@
Then, according to [5, Theorem A] and Remark Fake saddles and their transition maps 00footnotetext: 2020 AMS Subject Classification: 34D10, 34D20; Secondary: 34C05. 00footnotetext: Key words and phrases: singularities, Poincaré and Dulac transition maps, asymptotic expansion, uniform flatness, stability 00footnotetext: This work has been partially funded by the Ministry of Science, Innovation and Universities of Spain through the grant PID2021-125625NB-I00 and by the Agency for Management of University and Research Grants of Catalonia through the grant 2021SGR01015. , we obtain the following expression:
@@
To finish this work we apply the previous two results to the study of the following family of degenerated singularities considered in [3]:
@@
Notice that $f_{1}(x,0;\mu)\neq 0$ for all $x\in\mathbb{R}$ and $d(\mu)=4(1-c(\mu))-(a(\mu)-b(\mu))^{2}=\frac{2}{3}(4-\frac{1}{\beta})$ is positive if and only if $\beta>\frac{1}{4}$. By applying Theorem 2 we also deduce that the origin is monodromic if and only if $\beta>\frac{1}{4}$. In that case we can apply Theorem 4 and consider the limit (as $\eta\to 0^{+}$) transition map $\Pi_{-\infty}^{+\infty}(y;\mu)=e^{\gamma_{\pm}(\mu)}y+\mathcal{F}_{1+\epsilon}^{\infty}(\{\beta>\frac{1}{4}\})$ of $X_{\mu}$, with

### v3

@@ Abstract
1 Introduction and statements of main results
Following [6, 7] we define a fake saddle as a singular point having exactly two separatrices which are contained in a smooth invariant curve separating two hyperbolic sectors, see Figure 1. This kind of singularities are also known as impassable grains. Another way to think about a fake saddle is that the corresponding vector field can be put in a form that is like a degenerate flow box, i.e. near the singularity the phase portrait consists of parallel fibers, all but one of which have no singular points, and the singular fiber (the union of the singular point with its two separatrices) has a semi-stable equilibrium point, see [2]. Taking two transverse sections $\Sigma_{\alpha}$ and $\Sigma_{\omega}$ to the singular fiber outside the singular point there is a transition map $\Pi_{\alpha}^{\omega}:\Sigma_{\alpha}\to\Sigma_{\omega}$ which is well-defined on both sides of the singular fiber.
The local phase portrait of a singularity with non-zero linear part (for the nilpotent case see for instance [1, Theorem 3.5]) is well known and prevents the singular point from being a fake saddle. Notice that the smoothness of the singular fiber is strictly necessary because the Hamiltonian vector field $y\partial_{x}+x^{2}\partial_{y}$ has a nilpotent singularity at the origin with exactly two separatrices that are contained in the cuspidal curve $y^{2}-\frac{2}{3}x^{3}=0$ separating two hyperbolic sectors.
The first non-trivial and generic case of a fake saddle arises when the second-order jet is non-zero. Our objective is to characterize these generic fake saddles and to analyze their transition maps in order to determine whether the behavior is attractive or repulsive on each side of the singular fiber. The techniques we employ allow us to derive the leading term of an asymptotic expansion that is uniform with respect to parameters.
@@ 1 Introduction and statements of main results
This work is mainly motivated by the paper [2], which provides a normal form for generic fake saddles that we recall in the sequel. First we can locate the fake saddle at the origin $(0,0)$. Secondly, if the vector field is at least of class $\mathscr{C}^{3}$ there exists a local diffeomorphism that rectifies the germ of singular fiber, so that it is contained in the line $y=0$ in a certain coordinate system $(x,y)$, see [2, Lemma 2.1]. Thanks to the hypothesis that the second-order jet of the fake saddle is non-zero, after a suitable rescaling of the coordinates we can assume that the fake saddle is defined by a system of differential equations
@@
where $a,b,c\in\mathbb{R}$. This normal form motivates us to consider a smooth family of planar vector fields $\{X_{\mu}\}_{\mu\in\Omega\subset\mathbb{R}^{n}}$ having the form
@@
where $f_{1}(x,y;\mu)$, $f_{2}(x,y;\mu)$, $g_{1}(x,y;\mu)$ and $g_{2}(y;\mu)$ are $\mathscr{C}^{K}$-functions fulfilling $f_{1}(0,0;\mu)=f_{2}(0,0;\mu)=1$. We consider the following invariants
$a(\mu),\quad b(\mu):=g_{2}(0;\mu),\quad c(\mu):=g_{1}(0,0;\mu)$ (2)
of the family (1) and define the associated value $d(\mu):=4(1-c(\mu))-(a(\mu)-b(\mu))^{2}$, which will play a key role in the sequel.
@@
Theorem A.
If the invariants $(a,b,c)$ given in (2) of a vector field $X$ in the family (1) do not belong to $\{d=0\}\cap\{a^{2}-b^{2}=4\}$ then the origin is a fake saddle if and only if, either $d>0$, or $c=1$ and $a=b$. In both situations after blowing up the origin we have a single singular point on the exceptional divisor, which is a hyperbolic saddle of hyperbolicity ratio 1c>01-c>0 in the first one and a semi-hyperbolic saddle in the second one.
The hypothesis $(a,b,c)\notin\{d=0\}\cap\{a^{2}-b^{2}=4\}$ in Theorem A (which is always verified when $a=0$) can not be removed as Example 3 will show.
Let $\Omega$ be an open subset of $\mathbb{R}^{n}$ and let $W$ any subset of $\Omega$. We recall (cf. [4, Definition 1.2]) that a $\mathscr{C}^{K}$-function $f(s;\mu)$ defined in the intersection of $(0,+\infty)\times\Omega$ with an open neighborhood of $\{0\}\times\Omega\subset\mathbb{R}^{n+1}$ belongs to the flat class $\mathcal{F}_{L}^{K}(W)$ if for each $\mu_{0}\in W$ and $\nu=(\nu_{0},\ldots,\nu_{n})\in\mathbb{N}^{n+1}$ with $|\nu|=\nu_{0}+\cdots+\nu_{n}\leq K$ there exists $\varepsilon>0$ such that $|\partial_{s}^{\nu_{0}}\partial_{\mu_{1}}^{\nu_{1}}\cdots\partial_{\mu_{n}}^{\nu_{n}}f(s;\mu)|\leq Cs^{L-\nu_{0}}$ for every $s\in(0,\varepsilon)$ and $\mu\in\mathbb{B}_{\varepsilon}(\mu_{0})\cap\Omega$.
We consider the transverse sections $\Sigma_{\star}=\{x=\star\}$ to $y=0$, for $\star\in\{\alpha,\omega\}$, where $\alpha<0<\omega$. We assume that $f_{1}(x,0)>0$ for all $x\in[\alpha,\omega]$ and we consider the transition map $\Pi_{\alpha}^{\omega}:\Sigma_{\alpha}\to\Sigma_{\omega}$. Our second result extends [2, Theorem B], which only treats the case $a=b=0$, $\partial_{y}f_{1}=\partial_{y}g_{1}=0$, $f_{2}=1$ and $g_{2}=0$, when the transition map is restricted to $y>0$. In that theorem the authors characterize the stability of the transition map from the sign of the Cauchy principal value
$\mathrm{PV}\int_{\alpha}^{\omega}\frac{g_{1}(x,0;\mu)}{xf_{1}(x,0;\mu)}dx:=\lim\limits_{\varepsilon\to 0^{+}}\left(\int_{\alpha}^{-\varepsilon}\frac{g_{1}(x,0;\mu)}{xf_{1}(x,0;\mu)}dx+\int_{\varepsilon}^{\omega}\frac{g_{1}(x,0;\mu)}{xf_{1}(x,0;\mu)}dx\right).$
Theorem B.
Assume that $\mu_{0}\in\Omega$ and that the invariants (2) of the vector field $X_{\mu_{0}}$ in (1) belong to $\{d=4(1-c)-(a-b)^{2}>0\}$. If $\mu\approx\mu_{0}$ then the origin is a fake saddle of $X_{\mu}$ and its transition map $\Pi_{\alpha}^{\omega}:\Sigma_{\alpha}\to\Sigma_{\omega}$ satisfies $\Pi_{\alpha}^{\omega}(y;\mu)=e^{\gamma_{\pm}(\mu)}y+\mathcal{F}_{1+\epsilon}^{K}(\{d>0\})$ on $\pm y\geq 0$, where $\epsilon>0$ and
Using that $f_{1}(0,0)=1$ and $g_{1}(0,0)=c$ we can explicitly compute the previous Cauchy principal value using a convergent integral:
$\mathrm{PV}\int_{\alpha}^{\omega}\frac{g_{1}(x,0)}{xf_{1}(x,0)}dx=c\log\left|\frac{\omega}{\alpha}\right|+\int_{\alpha}^{\omega}\left(\frac{g_{1}(x,0)}{f_{1}(x,0)}-c\right)\frac{dx}{x}.$
$\square$
Clearly the sign of $\gamma_{\pm}(\mu)$ determines the stability of the singular fiber $y=0$ on the side $\pm y>0$. In the case $a=b=0$ the two values $\gamma_{\pm}(\mu)$ coincide with the principal value of the integral. Otherwise the stability of the two sides $\pm y>0$ can be different, see Example 3. $\square$
2 Proofs of the main results
Proof of Theorem A.
A straightforward computation shows that
@@ 2 Proofs of the main results
with $P(0,0)=1$ and $Q(0,v)=-v^{2}+(b-a)v+c-1$. The point $(u,v)=(0,0)$ is always a singular point of $Y$. If $d<0$ then $Y$ has two other singular points on the exceptional divisor $u=0$ with at least one non-zero eigenvalue. If $(a,b,c)\in\{d=0\}\setminus(\{c=1,a=b\}\cup\{a^{2}-b^{2}=4\})$ then $Y$ has another double singular point at (u,v)=(0,(ba)/2)(u,v)=(0,(b-a)/2) with a non-zero eigenvalue. In these situations the transition map $\Pi_{\alpha}^{\omega}:\{x=\alpha\}\to\{x=\omega\}$ is not well defined so that the origin is not a fake saddle of $X$. The remaining assertions are easy to check. ∎
Proof of Theorem B.
@@
Notice that $\lambda:=\lambda_{+}=1-c>0$ and and $\lambda_{-}=\frac{1}{\lambda_{+}}$. According to [5, p. 47] we consider the functions
@@
Then, according to [5, Theorem A] and Remark 1, we obtain the following expression:
@@
3 Examples and applications
For each integer $n\geq 3$ we consider the vector field
$X_{n}=(x+y)^{2}\partial_{x}+y^{n}\partial_{y}$
On the other hand, the blowup of the vector field $X_{4}$
Moreover, since $a=b=0$, according to Theorem B, the derivative at $v=0$ of the transition map $\Pi_{\alpha}^{\omega}(v)$ of $Y_{1}$ is
@@ 3 Examples and applications
To finish this work we apply Theorems A and B to the study of the following family of degenerated singularities considered in [3]:
@@
Notice that $f_{1}(x,0;\mu)\neq 0$ for all $x\in\mathbb{R}$ and $d(\mu)=4(1-c(\mu))-(a(\mu)-b(\mu))^{2}=\frac{2}{3}(4-\frac{1}{\beta})$ is positive if and only if $\beta>\frac{1}{4}$. By applying Theorem A we also deduce that the origin is monodromic if and only if $\beta>\frac{1}{4}$. In that case we can apply Theorem B and consider the limit (as $\eta\to 0^{+}$) transition map $\Pi_{-\infty}^{+\infty}(y;\mu)=e^{\gamma_{\pm}(\mu)}y+\mathcal{F}_{1+\epsilon}^{\infty}(\{\beta>\frac{1}{4}\})$ of $X_{\mu}$, with